OpenStudy (rissyroo118):

what will be the interest and maturity value of a 90-day note with a principal amount of $3,000 and 10% interest? $3,074 $3,000 $74 $2,926

OpenStudy (rissyroo118):

i think it is A) 3,000x0.010x90/365 =7.397260274+3,000= 3,007.397260274 rounds to being $3,074

OpenStudy (rissyroo118):

@Ismaelcarrillo

OpenStudy (rissyroo118):

@MrNood

OpenStudy (rissyroo118):

@mathmate

OpenStudy (scarlettfarra2000):

Have you tried solving this on your own?

OpenStudy (rissyroo118):

@scarlettFerra2000 yes i did it up above^^^

OpenStudy (mathmate):

"i think it is A) 3,000x0.010x90/365 =7.397260274+3,000= 3,007.397260274 rounds to being $3,074" The idea is good, but just missed the conversion of 10% to 0.010 instead of 0.10 as the correct value. So the working is: 3000*(1+0.10*90/365)=3000+73.973=3074 approximately.

OpenStudy (mathmate):

@ScarlettFarra2000 He did, in his first comment after posting the question.

OpenStudy (rissyroo118):

so instead of 0.010 i should put 0.10 , when i do i still get the same answer so i believe the correct is A @mathmate

OpenStudy (mathmate):

As I put earlier, $3074 is the correct answer. (note: I don't work with letters, because they are too easy to get swapped for the wrong one, and does not give the user a visual confirmation of the answer. You will need to do the translation to letters if that's how you write your answer.)